# Growth of Structural Lengthscale in Kob Andersen Binary Mixtures: Role of medium range order

### v1

@@ IV Results
In Fig.4(d), we now plot the $\Gamma_{bare}(r)$, which is the excess correlation between the bare SOP values of the A particles which are calculated at the single particle level, and find that as a function of temperature, it grows moderately, similar to that observed for all particles in Fig.1(c). However, when we plot the $\Gamma_{CG}(r)$, which is the excess correlation between the coarse-grained SOP values, we find that $\Gamma_{CG}(r)$ for the A particles at lower temperatures survives till longer distances. This clearly shows that the static lengthscale related to the coarse-grained order parameter grows. Note that this growth of excess correlation of the coarse-grained SOP at low temperatures can be visually seen in the form of patches of high and low SOP values in Fig.4(b). In Fig.4(f), we also plot the excess displacement-displacement correlation only for the $A$ particles. The plots for the excess correlation for the bare and coarse-grained SOP values and the excess displacement-displacement correlation for the B particles are plotted in Fig.12 in Appendix II.
@@
The static and dynamic lengthscales obtained for the $A$ and $B$ particles are shown in Fig.5. The static lengthscales obtained from $\Gamma_{CG}$ are also separately calculated for the A and the B particles by coarse-graining their SOP values till their respective $L_{max}$ values given in Fig.3(c). We find that the dynamic lengthscale of both species grows similarly. However, the static lengthscale of only the A particles grows like the dynamical lengthscale, but that of the B particles grows moderately, as shown in Fig.5.
@@
Interestingly, many of the earlier studies, trying to find the static correlation length for an 80:20 KALJ system, worked with LFS, like the 11A structure obtained from TCC Malins2013JCP , which is also the ⟨0,2,8⟩ Voronoï polyhedra Coslovich2011 . This structure was found to be long lived, thus contributing to the slow dynamics Royall2014 ; Coslovich2011 . Here we plot the distribution of the SOP of particles in the 11A structure obtained via TCC Malins2013JCP along with the distribution of the SOP of all the particles in Fig.6(a). We find that the SOP values of the 11A particles are shifted towards a higher caging potential, suggesting that these particles are indeed in a deeper caging potential. However, the distribution is also quite wide, which implies that, unlike proper crystals, the local arrangements of the particles in the 11A structure vary over a wide range. We plot the overlap function of the particles which are in the 11A structure along with the overlap function of all the particles in Fig.6(b). We find that, indeed, the particles belonging to the 11A structure are slower compared to the average dynamics. Note that at $T=0.45$ on average, there are about $12.5\%$ particles in the 11A structure. For a system of 16,000 particles, this comes to about 2000 particles. So we calculate the overlap function for the 2000 hardest particles and plot it in Fig.6(b). We find that they are marginally slower than the 11A particles. Thus, we show that our SOP can also identify slow moving particles. Our study reveals that the failure of the earlier studies in obtaining the growing static lengthscale in the 80:20 KALJ system is due to two possible reasons. The first is that the required coarse-graining to take care of the medium range order was not done, and the second is that many studies focused on LFS structures, which for this system is centred around the B particles that we know do not show much growth.
@@
We next study the growth of the static and dynamic lengthscale in a 60:40 mixture. This system is studied at a higher density $\rho=1.41$ so that the temperature range is similar to that of the 80:20 mixture (Fig.10 in Appendix I). In Fig.7(a) and Fig.7(b), we plot the excess displacement-displacement correlation functions of the A and the B particles, respectively. We find that similar to the 80:20 mixture, both for the A and B particles, the excess correlation grows with the lowering of the temperature. In Fig.8(a) and Fig.8(b), we plot the Spearman rank correlation of the coarse-grained SOP and the mobility at $t=\tau_{\alpha}$ for the A and B particles, respectively. Both show that there is an optimum value of coarse-graining length $L_{max}$ where the correlation is maximum, and $L_{max}$ grows with a decrease in temperature. Unlike in the case for the 80:20 mixture, in this case (Fig.8(c)), the $L_{max}$ for the $B$ particles grows almost as much as that for the $A$ particles.
@@
In Fig.8(d) and Fig.8(f), we show the snapshots of the bare SOP values for the A and the B particles, respectively, at $T=0.45$. These show that the values of the SOP fluctuate at the single particle level. In Fig.8(e) and Fig.8(g), we show snapshots of the coarse-grained SOP for the A and the B particles, respectively. We find that the coarse-grained SOP does show patches of regions with low and high SOP values. In Fig.8(h) and Fig.8(j), we plot $\Gamma_{bare}$ (Eq.17) at different temperatures, and they show a weak temperature dependence. However, $\Gamma_{CG}$ (Eq.22) for the A and the B particles show that with a decrease in temperature, the correlation survives till a longer distance suggesting a growing correlation length (Fig.8(i) and Fig.8(k)). To quantify these results in Fig.9 we plot the dynamic and the static lengthscales for the A and the B particles. We find that in this case, both the static and the dynamical lengthscales for both species grow together.
@@ V Conclusions
To examine the static counterpart, we analyse lengthscales derived from a recently developed structural order parameter (SOP) based on a mean field caging potential Manoj_prl_2021 ; Sharma2022 . When using the local bare SOP at a single particle level, the static lengthscale shows limited growth, echoing earlier observations Hocky2012 ; KARMAKAR20121001 ; Biroli2013 ; Berthier2016 ; Yaida2016 ; MalinsLifetimes2013 ; Royall2014 . This result is initially surprising given previous studies involving some of us, demonstrated a strong correlation between this SOP and dynamics Sharma2022 ; mohit_wca . To resolve this, we analyse the structure–dynamics correlation across different timescales. We find that while the local SOP correlates well with short time dynamics, its correlation weakens over long timescales. This is because long time rearrangements are influenced not only by the SOP of a given particle but also by the SOP of its neighbours, indicating the emergence of medium range order.
By coarse-graining the SOP over a spatial lengthscale, $L$, we identify an optimal coarse-graining length, $L_{max}$, that maximises structure–dynamics correlation at $\tau_{\alpha}$ time scale. We find that this $L_{max}$ grows with a decrease in temperature. For the 80:20 mixture, this growth is more pronounced for the A particles than for the B particles, while for the 60:40 mixture, both species show comparable growth. The corresponding static correlation length, extracted from the spatial correlations of the coarse-grained SOP, also increases significantly at low temperatures. Moreover, we find that for the 80:20 mixture, only the static lengthscale of A particles grows with cooling, and in the 60:40 mixture, the static lengthscales of both species grow in tandem with the dynamical lengthscale. At this point we do not completely understand why the structural lengthscale for the B particles do not grow but as discussed later there might be some correlation with the underlying crystalline order.
@@
Finally, we investigate whether the observed structural correlations are linked to underlying crystalline order. Previous analyses involving some of us showed that in the 80:20 mixture, the A particles exhibit a tendency to form fcc structures, while no evidence of CsCl-like ordering was observed Ujjwal_jcp_2016 . This indicates that B particles do not participate in the crystallization process in this composition. In contrast, the 60:40 mixture displays a tendency to form a CsCl structure, which involves both A and B particles Ujjwal_jcp_2016 . Interestingly, neither composition exhibits locally favored structures that correspond to the underlying crystalline order. Nevertheless, we find that the static length scale grows predominantly for the particle species associated with the dominant crystallization tendency, namely, the $A$ particles in the 80:20 mixture, and both $A$ and $B$ particles in the 60:40 mixture. This suggests a potential link between the growth of structural correlations and latent crystallization pathways. However, further investigations are needed to establish a definitive connection.
@@
Appendix I: Dynamics in 80:20 and 60:40 composition in KALJ systems
Figure 10: The $\alpha$-relaxation time obtained from the overlap function (Eq.2), $\tau_{\alpha}$, for two systems: 80:20 composition at $\rho=1.2$ and 60:40 composition at $\rho=1.41$ as a function of temperature. The solid lines are guides to the eye.
@@
Appendix II: Lengthscale studies of B-type particles in 80:20 composition
@@
In Fig.11 we plot the Spearman rank correlation $C_{R}(\overline{\beta\phi},\mu)$ against time scale by $\tau_{\alpha}$ which quantifies the relationship between the coarse-grained SOP and particle mobility for $B$ particles. As shown in Fig.11(a), for $L=0$ (i.e., the bare SOP), the correlation is initially strong at short times but decays rapidly, indicating that the bare SOP captures structural features relevant only to short time dynamics. This decay becomes more pronounced at lower temperatures, as illustrated in Fig.11(b). To predict dynamics over longer timescales, it becomes necessary to incorporate medium range structural information. This is evident from the increased structure–dynamics correlation when the SOP is coarse-grained over a spatial length $L$; up to a certain optimum value, $L_{max}$, the coarse-grained SOP retains a stronger correlation with mobility at longer times. Here, in Fig.12, we show the snapshots of the structural and dynamical parameters and their respective excess correlations for the B particles. We find that even for the coarse-grained SOP, the size of the patches representing high and low SOP values is smaller than that of the A particles (see Fig.4(b)). This leads to the smaller values of $L_{max}$ and $\xi_{CG}$ for the B particles compared to the A particles.
@@
Figure 11: Spearman rank correlation, $(C_{R}(\overline{\beta\Phi},\mu))$, between the coarse-grained SOP, $\overline{\beta\phi}$, and mobility, $\mu$, calculated at times scaled by the $\alpha$-relaxation time, $t/\tau_{\alpha}$, for B-type particles at different coarse-graining lengths, $L$. (a) Result at high temperature, $T=0.70$ and (b) at low temperature, $T=0.45$. The colour coding in (b) matches that of (a).
@@
Figure 12: Snapshots of structural and dynamical properties and their corresponding excess correlations for B-type of particles for 80:20 KALJ system at T=0.45. (a) Snapshot of the bare SOP, $\beta\phi$, appears spatially noisy, lacking visible large scale structure. (b)The excess correlation of the bare SOP, $\Gamma_{\text{bare}}(r)$, (Eq.17) confirms the absence of growth in bare static lengthscale.(c) Snapshot of the coarse-grained SOP, $\overline{\beta\phi}$, with $L_{max}$, shows emergence of some small patches of high and low SOP values. However these patches are much smaller than that observed for the A-type of particles (see Fig.4(b)) (d)The excess correlation of the coarse-grained SOP, $\Gamma_{\text{CG}}(r)$, (Eq.22) shows an increasing correlation length with decreasing temperature. Consistent with the size of the patches the static lengthscales obtained from the excess correlation are smaller than that of the A-type of particles (see Fig.5) (e)Snapshot of the particle mobility field, $\mu(\tau_{\alpha})$, calculated using Eq.19 over the $\alpha$-relaxation time shows low and high mobility domains. (f) The excess displacement–displacement correlation function, $\Gamma_{uu}(r,t_{\alpha})$, (Eq.7) exhibits growth with decreasing temperature.The colour coding in (d) and (f) matches with (b).

### v2

@@ IV Results
For each iso configurational run $\Gamma_{CG}(r)$ and the corresponding $\xi_{CG}$ is calculated by using the corresponding $L_{max}$ value. In Fig.4(d), we now plot the $\Gamma_{bare}(r)$, which is the excess correlation between the bare SOP values of the A particles, which are calculated at the single particle level, and find that as a function of temperature, it grows moderately. However, when we plot the configuration averaged $\Gamma_{CG}(r)$, which is the excess correlation between the coarse-grained SOP values, we find that $\Gamma_{CG}(r)$ for the A particles at lower temperatures survives till longer distances. This clearly shows that the static lengthscale related to the coarse-grained order parameter grows. Note that this growth of excess correlation of the coarse-grained SOP at low temperatures can be visually seen in the form of patches of high and low SOP values in Fig.4(b). In Fig.4(f), we also plot the excess displacement-displacement correlation only for the $A$ particles. The plots for the excess correlation for the bare and coarse-grained SOP values and the excess displacement-displacement correlation for the B particles are plotted in Fig.15 in Appendix IV.
@@
The static lengthscales obtained from $\Gamma_{CG}$ for every configuration are also separately calculated for the A and the B particles by coarse-graining their SOP values till their respective $L_{max}$ values obtained for that configuration. The average of this static lengthscale with error bars is plotted in Fig.5 along with the dynamic lengthscales. We find that the dynamic lengthscale of both species grows similarly. However, the static lengthscale of only the A particles grows like the dynamical lengthscale, but that of the B particles grows moderately, as shown in Fig.5. In Appendix IV in Fig.13, plot the $L_{max}$ against the $\xi_{CG}$ of the $A$ particles, and we find that they show a near linear dependence, suggesting that $\xi_{CG}$ grows in tandem with $L_{max}$.
In Fig.12 in Appendix III we show that the relaxation time can be fitted to a Vogel-Fulcher-Tammann (VFT) expression, $\tau_{\alpha}=\tau_{o}*exp(\frac{T_{VFT}}{K_{VFT}(T-T_{VFT})})$ where $\tau_{o}$ is the prefactor, $K_{VFT}$ is the fragility which quantifies how fast the dynamics diverges and $T_{VFT}$ is the temperature where the dynamics is expected to diverge. Similarly, the temperature dependence of the lengthscale is also given by the following expression $\xi=\xi_{o}[(T-T_{VFT})/T_{VFT}]^{-2/3}$, where $\xi_{o}$ is a prefactor. In Fig.5 we show that this functional form can reasonably fit the static and the dynamic lengthscales. This implies that there is a power law dependence between the relaxation time and the correlation length, where $ln(\tau_{\alpha})$ grows as $\xi^{d/2}$, shown earlier in other studies, where ‘d’ is the dimension tanakamrco_critical_natmat_2010 ; tanakamrco_kawasaki_prl_2007 ; TongTanaka2018 . In the inset of Fig.5 we plot $ln(\tau_{\alpha})$ against $\xi^{3/2}$, which can be fitted to a straight line, confirming the relationship. This temperature dependence of the lengthscales and its correlation with the dynamics is consistent with Ising-type critical scenario tanakamrco_critical_natmat_2010 ; Tanaka2012bondliquid ; Langer2013 and random first order transition theory in finite dimensionsLubchenko2007 ; Kirkpatrick1989 .
@@
Interestingly, many of the earlier studies, trying to find the static correlation length for an 80:20 KALJ system, worked with LFS, like the 11A structure obtained from TCC Malins2013JCP , which is also the ⟨0,2,8⟩ Voronoï polyhedra Coslovich2011 . This structure was found to be long lived, thus contributing to the slow dynamics Royall2014 ; Coslovich2011 . Here we plot the distribution of the SOP of particles in the 11A structure obtained via TCC Malins2013JCP along with the distribution of the SOP of all the particles in Fig.6(a). We find that the SOP values of the 11A particles are shifted towards a higher caging potential, suggesting that these particles are indeed in a deeper caging potential. However, the distribution is also quite wide, which implies that, unlike proper crystals, the local arrangements of the particles in the 11A structure vary over a wide range. We plot the overlap function of the particles that are in the 11A structure, along with the overlap function of all the particles in Fig.6(b). We find that, indeed, the particles belonging to the 11A structure are slower compared to the average dynamics. Note that at $T=0.45$ on average, there are about $12.5\%$ particles in the 11A structure. For a system of 16,000 particles, this comes to about 2000 particles. So we calculate the overlap function for the 2000 particles which have the deepest caging potential and plot it in Fig.6(b). We find that they are marginally slower than the 11A particles. Thus, we show that our SOP can also identify slow moving particles, and like the recently observed anticorrelation between the LFS and the excitations, there will be an anticorrelation between the particle in the deepest minima and the excitations Royall_excitations_2025 .
We next discuss the importance of coarse-graining the order parameter to obtain a proper lengthscale. Our results show that a single particle structural property can capture dynamical behaviour only at short timescales. In contrast, the dynamics and associated properties at the $\alpha$ relaxation timescale depend not only on the immediate structural environment of a tagged particle but also on that of its neighbours. For example, a particle in a shallow caging potential may appear mobile at short times, leading to a strong correlation between the structural order parameter and short-time dynamics. However, for the particle to remain mobile over longer times, its neighbours must also reside in shallow cages and thus be more mobile. This collective effect of the surrounding environment is naturally incorporated in a coarse-grained SOP, which, by including both the particle’s structure and that of its neighbours, provides the essential information needed to describe long-time dynamics. It is instructive to compare our findings with the earlier study of Zhang and Kob Zhang2020 , who correlated a structural order parameter with dynamics at the shorter timescale corresponding to the peak of the non-Gaussian parameter, $t^{*}$. The difference between $t^{*}$ and the structural relaxation time $\tau_{\alpha}$ increases as temperature decreases, with $\tau_{\alpha}$ becoming much larger. Consequently, their order parameter, like our bare SOP, showed little growth at lower temperatures. Our study suggests that this slow growth arises because an order parameter that successfully predicts short time dynamics is not necessarily appropriate for describing dynamics and dynamical properties at longer times.
@@
We next study the growth of the static and dynamic lengthscale in a 60:40 mixture. This system is studied at a higher density $\rho=1.41$ so that the temperature range is similar to that of the 80:20 mixture (Fig.12 in Appendix III). In Fig.7(a) and Fig.7(b), we plot the excess displacement-displacement correlation functions of the A and the B particles, respectively. We find that, similar to the 80:20 mixture, both for the A and B particles, the excess correlation grows with the lowering of the temperature. In Fig.8(a) and Fig.8(b), we plot the Spearman rank correlation of the coarse-grained SOP and the mobility at $t=\tau_{\alpha}$ for the A and B particles, respectively. Both show that there is an optimum value of coarse-graining length $L_{max}$ where the correlation is maximum, and $L_{max}$ grows with a decrease in temperature. Unlike in the case for the 80:20 mixture, in this case (Fig.8(c)), the $L_{max}$ for the $B$ particles grows almost as much as that for the $A$ particles.
@@
In Fig.8(d) and Fig.8(f), we show the snapshots of the bare SOP values for the A and the B particles, respectively, at $T=0.45$. These show that the values of the SOP fluctuate at the single particle level. In Fig.8(e) and Fig.8(g), we show snapshots of the coarse-grained SOP for the A and the B particles, respectively. We find that the coarse-grained SOP does show patches of regions with low and high SOP values. In Fig.8(h) and Fig.8(j), we plot $\Gamma_{bare}$ (Eq.17) at different temperatures, and they show a weak temperature dependence. However, $\Gamma_{CG}$ (Eq.22) for the A and the B particles show that with a decrease in temperature, the correlation survives till a longer distance, suggesting a growing correlation length (Fig.8(i) and Fig.8(k)). To quantify these results in Fig.9, we plot the dynamic and the static lengthscales for the A and the B particles. We find that in this case, both the static and the dynamical lengthscales for both species grow together. Moreover, similar to the 80:20 mixture, we show that even for the 60:40 mixture, the expression $\xi=\xi_{o}[(T-T_{VFT})/T_{VFT}]^{-2/3}$ can fit the static and the dynamic lengthscales reasonably well. In the inset, we also plot $ln(\tau_{\alpha})$ against $\xi^{3/2}$, which confirms the power law relationship between the relaxation time and the lengthscales.
@@ V Conclusions
To examine the static counterpart, we analyse lengthscales derived from a recently developed structural order parameter (SOP) based on a mean field caging potential Manoj_prl_2021 ; Sharma2022 . When using the local bare SOP at a single particle level, the static lengthscale shows limited growth, echoing earlier observations Hocky2012 ; KARMAKAR20121001 ; Biroli2013 ; Berthier2016 ; Yaida2016 ; MalinsLifetimes2013 ; Royall2014 . This result is initially surprising given previous studies involving some of us demonstrated a strong correlation between this SOP and dynamics Sharma2022 ; mohit_wca . To resolve this, we analyse the structure–dynamics correlation across different timescales. We find that while the local SOP correlates well with short time dynamics, its correlation weakens over long timescales. This is because long time rearrangements are influenced not only by the SOP of a given particle but also by the SOP of its neighbours, indicating the emergence of medium range order.
By coarse-graining the SOP over a spatial lengthscale, $L$, we identify an optimal coarse-graining length, $L_{max}$, that maximises structure–dynamics correlation at $\tau_{\alpha}$ time scale. We find that this $L_{max}$ grows with a decrease in temperature. For the 80:20 mixture, this growth is more pronounced for the A particles than for the B particles, while for the 60:40 mixture, both species show comparable growth. The corresponding static correlation length, extracted from the spatial correlations of the coarse-grained SOP, also increases significantly at low temperatures. Moreover, we find that for the 80:20 mixture, only the static lengthscale of A particles grows with cooling, and in the 60:40 mixture, the static lengthscales of both species grow in tandem with the dynamical lengthscale. At this point, we do not completely understand why the structural lengthscale for the B particles does not grow, but as discussed later, there might be some correlation with the underlying crystalline order.
@@
In recent work, Tong and Tanaka have made an important and insightful observation regarding the common structural packing in systems with and without medium range crystalline order TongTanaka2018 ; tanakamrco_kawasaki_JPCM_2010 . It is well known that monodisperse or weakly polydisperse systems in two dimensions exhibit hexatic ordering of particles, which is directly related to the corresponding crystal structure. However, many supercooled liquid systems do not show any such ordering. Tong and Tanaka demonstrated that even in these systems, there exists a form of local ordering corresponding to the most sterically favored structure. Although this local order is not connected to any crystalline symmetry, it can be captured by their newly proposed order parameters, $\theta$ in two dimensions and $\Omega$ in three dimensions. This finding suggests that, for both systems with and without MRCO, the underlying tendency of a supercooled liquid is to achieve the most efficient local packing. In our study, the structural order parameter essentially characterizes the cage formed by the first-neighbor shell. This description naturally encodes information about local packing and is therefore similar in spirit to the order parameters introduced by Tong and Tanaka TongTanaka2018 ; TongTanaka2020 . Interestingly, and consistent with our present observations, it has been shown earlier that coarse-graining of $\theta$ and $\Omega$ successfully captures the growth of a static lengthscale even in MRCO-free systems, both in two and three dimensions TongTanaka2018 ; TongTanaka2020 .
Finally, we investigate whether the observed structural correlations are linked to underlying crystalline order. Previous analyses involving some of us showed that in the 80:20 mixture, the A particles exhibit a tendency to form fcc structures, while no evidence of CsCl-like ordering was observed Ujjwal_jcp_2016 . This indicates that B particles do not participate in the crystallisation process in this composition. In contrast, the 60:40 mixture displays a tendency to form a CsCl structure, which involves both A and B particles Ujjwal_jcp_2016 . Although in our simulations we do not see a crystallisation of the 60:40 mixture, we know that it is very close to the composition which crystallises, i.e., the 58:42 mixture, as shown in Appendix V, crystallises primarily to a CsCl structure. Interestingly, neither the 80:20 nor the 60:40 composition exhibits growing locally favoured structures that correspond to the underlying crystalline order. Nevertheless, we find that the static lengthscale grows predominantly for the particle species associated with the dominant crystallisation tendency, namely, the $A$ particles in the 80:20 mixture, and both $A$ and $B$ particles in the 60:40 mixture. This suggests a potential link between the growth of structural correlations and latent crystallisation pathways. However, further investigations are needed to establish a definitive connection.
@@
Appendix I: Temperature dependence of the distribution of the local SOP, $\beta\phi$
The depth of the mean field caging potential, for a particle, $\beta\phi$, is derived under the mean field approximation, which assumes a frozen background by neglecting the dynamics of the surrounding particles. We can compute it for every particle by using the information of the local rdf and the distribution $P(\beta\phi)$ directly reflects spatial heterogeneity. With decreasing temperature, $P(\beta\phi)$ shifts to larger values as shown in Fig. 10, indicating more particles in deeper cages. Thus $\beta\phi$ captures the structural heterogeneity and its temperature evolution.
Figure 10: The temperature dependence of the distribution of the bare SOP, $\beta\phi$, for the 80:20 composition.
Appendix II: Reliability of O.Z. Fitting in $S_{4}(q,t)$
Figure 11: Scaling plot of $S_{4}(q,\tau_{\alpha})/S_{4}(q{=}0,\tau_{\alpha})$ versus $q\xi_{4}$. The dotted line represents the Ornstein-Zernike function used for fitting.
In Fig. 11, we show a scaling plot of $S_{4}(q,t)$ together with the Ornstein-Zernike (OZ) function (dotted line). The data exhibit good overlap with the OZ function at all temperatures, demonstrating that the fitting is reliable Szamel_s4qt ; Tah_dis_dis .
Appendix III: Dynamics in 80:20 and 60:40 composition in KALJ systems
Figure 12: The $\alpha$-relaxation time obtained from the overlap function (Eq.2), $\tau_{\alpha}$, for two systems: 80:20 composition at $\rho=1.2$ and 60:40 composition at $\rho=1.41$ as a function of temperature. The solid lines are VFT fits, $\tau_{\alpha}=\tau_{o}*exp(\frac{T_{VFT}}{K_{VFT}(T-T_{VFT})})$ where $\tau_{o}$ is the prefactor, $K_{VFT}$ is the fragility which quantifies how fast the dynamics diverges, and $T_{VFT}$ is the temperature where the dynamics is expected to diverge. For 80:20 mixture $T_{VFT}=0.319$ and for 60:40 mixture $T_{VFT}=0.312$
@@
Appendix IV: Lengthscale studies in 80:20 composition
Figure 13: In 80:20 mixture for A particles, the plot of the structural correlation length, $\xi_{\mathrm{CG}}$, versus the optimal coarse-graining lengthscale, $L_{\max}$, across all studied temperatures. The two measures exhibit a strong positive, near-linear relationship, indicating that $\xi_{\mathrm{CG}}$ grows in tandem with the $L_{\max}$.
In Fig.13 we plot the structural correlation length, $\xi_{CG}$ against $L_{max}$, which shows that the former grows in tandem with the latter.
@@
In Fig.14 we plot the Spearman rank correlation $C_{R}(\overline{\beta\phi},\mu)$ against the time scale by $\tau_{\alpha}$, which quantifies the relationship between the coarse-grained SOP and particle mobility for $B$ particles. As shown in Fig.14(a), for $L=0$ (i.e., the bare SOP), the correlation is initially strong at short times but decays rapidly, indicating that the bare SOP captures structural features relevant only to short time dynamics. This decay becomes more pronounced at lower temperatures, as illustrated in Fig.14(b). To predict dynamics over longer timescales, it becomes necessary to incorporate medium range structural information. This is evident from the increased structure–dynamics correlation when the SOP is coarse-grained over a spatial length $L$; up to a certain optimum value, $L_{max}$, the coarse-grained SOP retains a stronger correlation with mobility at longer times. Here, in Fig.15, we show the snapshots of the structural and dynamical parameters and their respective excess correlations for the B particles. We find that even for the coarse-grained SOP, the size of the patches representing high and low SOP values is smaller than that of the A particles (see Fig.4(b)). This leads to the smaller values of $L_{max}$ and $\xi_{CG}$ for the B particles compared to the A particles.
@@
Figure 14: Spearman rank correlation, $(C_{R}(\overline{\beta\Phi},\mu))$, between the coarse-grained SOP, $\overline{\beta\phi}$, and mobility, $\mu$, calculated at times scaled by the $\alpha$-relaxation time, $t/\tau_{\alpha}$, for B-type particles at different coarse-graining lengths, $L$. (a) Result at high temperature, $T=0.70$ and (b) at low temperature, $T=0.47$. The colour coding in (b) matches that of (a).
@@
Figure 15: Snapshots of structural and dynamical properties and their corresponding excess correlations for B-type particles for 80:20 KALJ system at T=0.45. (a) Snapshot of the bare SOP, $\beta\phi$, appears spatially noisy, lacking visible large scale structure. (b)The excess correlation of the bare SOP, $\Gamma_{\text{bare}}(r)$, (Eq.17) confirms the absence of growth in bare static lengthscale.(c) Snapshot of the coarse-grained SOP, $\overline{\beta\phi}$, with $L_{max}$, shows emergence of some small patches of high and low SOP values. However these patches are much smaller than that observed for the A-type of particles (see Fig.4(b)) (d)The excess correlation of the coarse-grained SOP, $\Gamma_{\text{CG}}(r)$, (Eq.22) shows an increasing correlation length with decreasing temperature. Consistent with the size of the patches the static lengthscales obtained from the excess correlation are smaller than that of the A-type of particles (see Fig.5) (e)Snapshot of the particle mobility field, $\mu(\tau_{\alpha})$, calculated using Eq.19 over the $\alpha$-relaxation time shows low and high mobility domains. (f) The excess displacement–displacement correlation function, $\Gamma_{uu}(r,t_{\alpha})$, (Eq.8) exhibits growth with decreasing temperature.The colour coding in (d) and (f) matches that of (b).
Appendix V: Crystallisation in KALJ 58:42 composition
Figure 16: (a) Potential energy, P.E., as a function of time, t, at $T=0.66$, showing a sudden drop indicative of a crystallisation event. We mark two regions before and after the drop as pre-crystalisation and post-crystalisation. (b) Probability distribution of the averaged bond orientational order parameter Steinhardt1983 ; Mickel2013_boo $\overline{q}_{6}$(Eq.24) for B-type particles with respect to A-type neighbours in the two regions i.e pre-crystalisation and post-crystalisation. A clear shift is observed from the pre-crystallisation state (black, circles) to the post-crystallisation state (red, squares), with the distribution moving to a $\bar{q}_{6}$ value associated with bcc like local order. For comparison, the $\overline{q}_{6}$ distribution for the 50:50 KALJ composition (green, triangle) which is known to crystallize into a CsCl structure Atreyee_crystal_jcp is also plotted.
In the 58:42 composition of the KALJ system, we observe clear signatures of crystallisation at a temperature $T=0.66$. As shown in Fig. 16(a), the time evolution of the system’s potential energy exhibits a sudden drop, consistent with a crystallisation event. To further characterise this transition, we analysed the local bond orientational order using the averaged $\overline{q}_{l}$ parameter for $l=6$ given by Steinhardt1983 ; Mickel2013_boo ; Ujjwal_jcp_2016 :
$\bar{q}_{l}=\sqrt{\frac{4\pi}{2l+1}\sum_{m=-l}^{l}\left|\bar{q}_{lm}\right|^{2}}$ (24)
where, $\bar{q}_{lm}(i)=\frac{1}{{N}_{i}}\sum_{k=0}^{{N}_{i}}q_{lm}(k)$ and $q_{lm}(i)=\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}Y_{lm}\left(\theta(r_{ij}),\phi(r_{ij})\right)$. Here, $Y_{lm}$ are spherical harmonics, and $N_{i}$ is the number of neighbors of particle $i$ within a cutoff radius $r_{\text{cut}}=1.2$. The quantity $\bar{q}_{6}$ characterises the degree of six-fold local orientational symmetry, averaged over the neighbourhood of particle $i$. In Fig. 16(b), we present the distribution of $\overline{q}_{6}$ values computed for B-type particles with respect to their neighbouring A-type particles. Notably, after the energy drop in Fig.16(b), the distribution of BOO shifts to the range which is characteristic of bcc symmetry ($\bar{q_{6}}=0.6$). This behaviour closely correlates with the local structural environment found in the well-established CsCl-type crystalline phase, typically observed in the equimolar (50:50) KALJ composition at low temperatures Atreyee_crystal_jcp . The similarity in the $\overline{q}_{6}$ distributions between the 58:42 and 50:50 compositions further supports the presence of CsCl ordering in the 58:42 KALJ system.